OpenStudy (anonymous):

Can anyone help me verify the identity of: (cscθ + cotθ)/ tanθ + sinθ = cscθcotθ Would appreciate it.

OpenStudy (anonymous):

I figured it out. It is a true identity. \[\cos \theta/ \sin^2 \theta = \cos \theta / \sin^2 \theta\]